OpenStudy (anonymous):

d/dx ( \int\limits_{2x}^{x^3} \1/(1+x^2)^(1/2)dt

OpenStudy (kira_yamato):

\[\frac{d}{dx} \int\limits_{2x}^{x^3} \frac{1}{\sqrt{(1+x^2)}}dt = \frac{d}{dx} \frac{t}{\sqrt{(1+x^2)}} = \frac{d}{dx}\frac{x^3 - 2x}{\sqrt{1 + x^2}}\]

OpenStudy (kira_yamato):

You can do quotient rule from there